### v1

# Dichotomy theorem distinguishing non-integrability and the lowest-order Yang-Baxter equation for isotropic spin chains

Abstract
We investigate the integrability and non-integrability of isotropic spin chains with nearest-neighbor interaction with general spin $S$. We prove a dichotomy theorem that a single relation sharply separates two scenarios: (i) this system is non-integrable, or (ii) the lowest-order Yang-Baxter equation is satisfied. This result solves in the affirmative the Grabowski-Mathieu conjecture stating that a model is integrable only if this model has a 3-local conserved quantity. This theorem also serves as a complete classification of integrability and non-integrability for $S\leq 13.5$, suggesting that all the integrable models are in the scope of the Yang-Baxter equation.
Introduction.— Integrability and non-integrability of quantum many-body systems are important subjects in theoretical physics. Integrable systems have been studied for almost 100 years, and various quantum systems such as the Heisenberg model, the transverse Ising model, and more complicated models were discovered as integrable models JM ; Bax . We note that unlike classical integrability there is no universally-accepted definition of quantum integrability and non-integrability, though known integrable models with local interactions share common properties including solvability and infinitely many local conserved quantities. Following previous studies CM11 ; GE16 ; Shi19 , in this Letter, we adopt a working definition of integrability and non-integrability as having infinitely many and no nontrivial local conserved quantities, respectively.
To find and solve integrable models, a sophisticated technique based on the Yang-Baxter equation, named the algebraic Bethe ansatz or the quantum inverse scattering methods, works as a powerful tool Fad ; KBI . Precisely, if the Yang-Baxter equation is satisfied, then we can obtain an infinite series of local conserved quantities DG ; FF ; GM94 ; GM95-1 ; NF20 , with which energy eigenstates can be constructed. In this case, we also have a boost operator, which generates local conserved quantities through a simple calculation GM94 ; GM95-1 . It is noteworthy, however, that some important integrable models including the Hubbard model and the XY model neither satisfy the conventional (differentiable) Yang-Baxter equation nor have a boost operator GM95-2 ; Lin01 ; Lee21 .
In contrast to integrability, theoretical aspects of non-integrability have not yet been fully addressed, though non-integrability (absence of nontrivial local conserved quantities) is a relevant property to various phenomena including thermalization Lan ; EF ; SM17 ; MS17 ; SMreply and response to external drives Kub ; SK ; SF ; Zot . One reason for this insufficiency is that the proof of non-existence (no solution or no nontrivial local conserved quantity) is usually harder than the proof of existence (of solutions and nontrivial local conserved quantities). Despite the scarcity of studies focusing on non-integrability, the existing literature can broadly be classified into two approaches.
The first approach to non-integrability is a traditional one, examining solvability and unsolvability by restricting the solution method to the Yang-Baxter equation KS82 ; Ken92 ; BY94 ; BN14 ; Hie92 ; LPR19 ; Lee20 ; Lee21 ; Mai24 ; LP24 . With this approach, several classes of spin chains have been classified into solvable and unsolvable ones in terms of the Yang-Baxter equation. As far as previous research has explored, all models satisfying the lowest-order Yang-Baxter equation (the Reshetikin condition KS82 ; Ken92 ) also satisfy the full-order Yang-Baxter equation, implying the solvability. However, this approach suffers from an inherent limitation: we cannot rule out the possibility that a model is not solvable by the Yang-Baxter equation but is still solvable by another, possibly yet unknown, solution method.
The second approach possesses generality, where the presence or absence of local conserved quantities is directly examined. A pioneering investigation in this direction was carried out by Grabowski and Mathieu GM95-2 , where they put two conjectures on spin chains with nearest-neighbor interaction: (1) A model is integrable only if it has a 3-local conserved quantity. (2) A model is integrable if it satisfies the lowest-order Yang-Baxter equation. Conjecture 1 directly concerns non-integrability, while conjecture 2 is closely aligned with the field of integrable systems, which is related to the first approach. Henceforth, we refer to conjecture 1 as the Grabowski-Mathieu conjecture otherwise explicitly noted. The Grabowski-Mathieu conjecture and its extension GP21 , if proved, are expected to show its significance when it is applied to previously unexplored systems, since it serves as the guidepost of investigation.
Recently, a mathematical technique to prove non-integrability was invented Shi19 , and by applying this technique various quantum spin systems have been proven to be non-integrable Chi24-1 ; Shi24 ; PL24-1 ; CY24 ; ST24 ; Chi24-2 ; PL24-2 ; HYC24 ; YCS24-1 ; YCS24-2 ; Shi25 ; FT25 . Notably, symmetric $S=1/2$ systems in one YCS24-1 ; YCS24-2 and higher dimensions ST24 ; Chi24-2 are completely classified into integrable and non-integrable, which confirms the Grabowski-Mathieu conjecture in the affirmative within this class. Another important progress is brought by Hokkyo Hok25 , which observes similarity among many previous non-integrability proofs and clarifies the sufficient conditions for proving non-integrability. He showed that if two operator equations respectively have solutions only in specific forms, then the same non-integrability proof works. However, these classification theorems are derived by exhaustively proving the integrability and non-integrability of all models in this class, which differs from the original aim of the Grabowski-Mathieu conjecture as the guidepost for investigation of wide classes of systems that include many unexplored systems. An exceptional general statement Hok25 employs not simple criteria, and thus what concrete systems adopt these results are still not clear.
In this Letter, we bridge the existing two approaches and establish a dichotomy theorem for isotropic spin chains with general spin $S$. Precisely, we prove that whether a single equation contains a certain type of term or not completely determines whether this model is non-integrable or satisfies the lowest-order Yang-Baxter equation. As its direct consequence, any integrable system is shown to have 3-local and 4-local conserved quantities generated by a boost operator, and hence the Grabowski-Mathieu conjecture is solved in the affirmative for isotropic spin chains. Importantly, the separation into the two extreme scenarios applies all isotropic spin chains including unexplored ones, which indeed meets the aim of the Grabowski-Mathieu conjecture. We also note that our criterion is tested by a simple and direct calculation, which shows clear contrast to existing results GM95-2 ; Hok25 requiring to solve search problems of operator equations.
Setup and main result.— We consider a shift-invariant isotropic spin chain with general spin $S$ with nearest-neighbor interaction. The isotropic condition is equivalent to the $SU(2)$ invariance. The system Hamiltonian in consideration is expressed as
$H=$ $\sum_{i}h_{i,i+1},\hskip 15.0pth_{i,i+1}=\sum_{n=1}^{2S}J_{n}(\bm{S}_{i}\cdot\bm{S}_{i+1})^{n}.$ (1)
Figure 1: Claim of our dichotomy theorem. If $D$ has no 3-support operator, then this model satisfies the lowest-order yang-Baxter equation (the Reshetikin condition) and has a boost operator generating 3-local and 4-local conserved quantities. If $D$ has a 3-support operator, then this model is non-integrable in the sense that this model has no nontrivial local conserved quantity.
Now we state the dichotomy theorem on the isotropic spin chains (see also Fig. 1).
Theorem.
Consider a shift-invariant isotropic spin chain with general spin $S$ with nearest-neighbor interaction, whose Hamiltonian is given by Eq. (1). For the above class of spin chains, whether the quantity
$D=[[h_{i,i+1},h_{i+1,i+2}],h_{i,i+1}+h_{i+1,i+2}]$ (2)
contains a nonzero 3-support operator or not completely determines whether the model is non-integrable or satisfies the lowest-order Yang-Baxter equation (the Reshetikin condition).
The latter case is equivalent to the existence of a boost operator in the form of $B=\sum_{j}jh_{j,j+1}$ generating 3-local and 4-local conserved quantities given by
$Q^{3}=\sum_{i}$ $[h_{i,i+1},h_{i+1,i+2}],$ (3)
$Q^{4}=\sum_{i}$ $2[[h_{i,i+1},h_{i+1,i+2}],h_{i+2,i+3}]$
$+[[h_{i,i+1},h_{i+1,i+2}],h_{i+1,i+2}].$ (4)
This boost operator is conjectured to generate higher conserved quantities $Q^{n}$ for $n\geq 5$ (conjecture 2 of Ref. GM95-2 ), while whether the generated operators are indeed conserved quantities is an open problem. The details of the lowest-order Yang-Baxter equation (the Reshetikin condition) and the boost operator are presented in SM .
This theorem reveals that the calculation of a single relation completely determines either non-integrability or the lowest-order Yang-Baxter equation. This shows a high contrast to the Grabowski-Mathieu conjecture GM95-2 and its partial proof by Hokkyo Hok25 , where we need to solve a search problem on operator equations. In addition, provided conjecture 2 of Grabowski and Mathieu GM95-2 , our theorem implies an astonishing claim that whether $D$ contains a nonzero 3-support operator or not completely determines the integrability and non-integrability of isotropic spin chains.
We here briefly comment on what happens if the above conjecture is false in isotropic spin chains. In this case, there exists an isotropic spin chain with a finite number of nontrivial local conserved quantities, which is called a partially-integrable system YCS24-1 ; YCS24-2 . Moreover, this counterexample is higher than or equal to $S=14$. These observations impress that conjecture 2 is highly plausible.
Implications of this theorem.— This theorem has various implications for integrability and non-integrability.
First, this theorem confirms that all integrable isotropic spin chains should have 3-local and 4-local conserved quantities in the form of Eqs. (3) and (4). Our theorem excludes the existence of (unknown) integrable systems without 3-local and 4-local conserved quantities. This proves the Grabowski-Mathieu conjecture GM95-2 in the affirmative in the class of isotropic spin chains, stating that a system is integrable only if it has a 3-local conserved quantity. This is the first proof of the Grabowski-Mathieu conjecture for a large class of spin systems including models which we have not yet explored.
In addition, our theorem fills the gap between the two approaches to non-integrability explained in the introduction. In particular, this dichotomy theorem justifies the validity of the first approach with fixing solution method Ken92 ; BY94 by confirming that the lowest-order Yang-Baxter equation does not miss integrable systems in isotropic spin chains. Given the fact that some integrable models including the Hubbard model do not satisfy the lowest-order Yang-Baxter equation, it is striking that all isotropic integrable models admit the lowest-order Yang-Baxter equation.
This theorem also serves as the complete classification of integrability and non-integrability up to $S=13.5$. Kennedy Ken92 and Batchelor and Yung BY94 have examined all isotropic spin chains satisfying the Reshetikin condition less than or equal to $S=13.5$ by computer search and found that all the spin chains satisfying the Reshetikin condition are in fact exhausted by known integrable systems solvable by the Yang-Baxter equation. Precisely, there are four infinite sequences of integrable systems and one exceptional integrable system for $S=3$ (see End matter), which correspond to four infinite series of compact Lie groups and one exceptional Lie group satisfying SU(2) symmetry BY94 . Combining this fact and our theorem, we conclude that these four infinite sequences of integrable systems and one exceptional integrable system are the complete list of integrable isotropic spin chains up to $S=13.5$.
Reshetikin condition and the Yang-Baxter equation.— Before going to the proof, we here clarify the relationship among conditions on integrable systems. The Yang-Baxter equation is a relation on a one-parameter family of a matrix $R_{ij}(\lambda)$ on two sites $i$ and $j$ given by
$R_{12}(\lambda)R_{23}(\lambda+\mu)R_{23}(\mu)=R_{23}(\mu)R_{13}(\lambda+\mu)R_{12}(\lambda).$ (5)
The $R$ matrix can be expanded as
$R(\lambda)=\Pi+\lambda\Pi h+\lambda^{2}\frac{1}{2}\Pi h^{2}+\cdots,$ (6)
where $h$ is the local Hamiltonian and $\Pi$ is a swap operator. The third and higher-order terms are not universal.
Plugging Eq. (6) into the Yang-Baxter equation (5) and compare the lowest nontrivial order, $\lambda\mu^{2}$, we obtain the Reshetikin condition claiming that $D$ in Eq. (2) can be expressed as KS82 ; Ken92
$D_{123}=E_{23}-E_{12},$ (7)
where $E_{ij}$ is a proper operator on two sites $i$ and $j$. Thus, the Reshetikin condition is the nontrivial lowest-order relation of the Yang-Baxter equation. The Reshetikin condition is also equivalent to the relation $[Q^{3},H]=0$ with $Q^{3}$ given by Eq. (3), meaning the presence of 3-local conserved quantity in this form. Note that the Reshetikin condition also implies $[Q^{4},H]=0$ with $Q^{4}$ given by Eq. (4) (see Supplemental Material SM ).
Useful basis.— To prove our main theorem, we employ a useful basis of operators on spin $S$ named noncommutative spherical harmonics, which is also known as fuzzy harmonics Iso01 ; LV14 ; AA19 . The basis operator $Y^{lm}$ has two indices, the azimuthal quantum number $l$ and the magnetic quantum number $m$, which take integers satisfying $0\leq l\leq 2S$ and $-l\leq m\leq l$. The product of two noncommutative spherical harmonics is given by
$Y^{l_{1},m_{1}}Y^{l_{2},m_{2}}=\sum_{l}K^{l_{1},l_{2},l}_{m_{1},m_{2}}Y^{l,m_{1}+m_{2}},$ (8)
where $K^{l_{1},l_{2},l}_{m_{1},m_{2}}$ is a coefficient proportional to the well-known Clebsch–Gordan coefficient. For our purpose, important properties of $K$ are the following two:
  • $K^{l_{1},l_{2},l}_{m_{1},m_{2}}$ is zero if the triangular condition $\left|l_{1}-l_{2}\right|\leq l\leq l_{1}+l_{2}$ is not satisfied.
  • $K^{l_{1},l_{2},l}_{m_{1},m_{2}}=K^{l_{2},l_{1},l}_{m_{2},m_{1}}$ if $l_{1}+l_{2}+l$ is even, and $K^{l_{1},l_{2},l}_{m_{1},m_{2}}=-K^{l_{2},l_{1},l}_{m_{2},m_{1}}$ if $l_{1}+l_{2}+l$ is odd.
The latter fact implies that a commutator $[Y^{l_{1},m_{1}},Y^{l_{2},m_{2}}]$ consists of $Y^{l^{\prime},m_{1}+m_{2}}$ such that $l^{\prime}$ and $l_{1}+l_{2}$ are different parities, and an anti-commutator $\{Y^{l_{1},m_{1}},Y^{l_{2},m_{2}}\}$ consists of $Y^{l^{\prime},m_{1}+m_{2}}$ such that $l^{\prime}$ and $l_{1}+l_{2}$ are the same parity. Using this basis, the local Hamiltonian $h$ given in Eq. (1) is also expanded as SM
$h_{i,i+1}=\sum_{l=0}^{2S}c_{l}\left(\sum_{m=-l}^{l}Y^{l,m}_{i}Y^{l,-m}_{i+1}\right),$ (9)
where we renormalized $Y^{lm}$ such that the coefficients of $Y^{l,m}_{i}Y^{l,-m}_{i+1}$ to be one in this expression.
Proof: former part.— We first consider the case that $D$ does not have 3-support operators with nonzero coefficients. By comparing our condition to the lowest-order Yang-Baxter equation (7), our task boils down to that 2-support operators in $D$ take the form of $E_{i+1,i+2}-E_{i,i+1}$. We shall show this fact as follows.
By assumption, $[[h_{i,i+1},h_{i+1,i+2}],h_{i+1,i+2}]$ is an (at most) 2-support operator, and since site $i$ has a nontrivial action, this 2-support operator is on $i$ and $i+1$, which we denote by
$[[h_{i,i+1},h_{i+1,i+2}],h_{i+1,i+2}]=A_{i,i+1}.$ (10)
Considering its reflection, we have
$[[h_{i,i+1},h_{i+1,i+2}],h_{i,i+1}]=$ $-[[h_{i+1,i+2},h_{i,i+1}],h_{i,i+1}]$
$=$ $-A_{i+2,i+1}.$ (11)
Thus, it suffices to show that $A$ has inversion symmetry;
$A_{ij}=A_{ji},$ (12)
which follows the argument below. Since both $[h_{i,i+1},h_{i+1,i+2}]$ and $h_{i+1,i+2}$ are $SU(2)$ symmetric, their commutator $[[h_{i,i+1},h_{i+1,i+2}],h_{i+1,i+2}]$ should also be $SU(2)$ symmetric. Notice that all 2-support $SU(2)$ symmetric operator takes the form of the right-hand side of Eq. (1), which also has inversion symmetry.
Proof: latter part.— We next consider the case that $D$ has a 3-support operator with nonzero coefficients. We prove the non-integrability of this Hamiltonian with the help of the scheme proposed by Hokkyo Hok25 . Below we denote the $m$-support part of $k$-local conserved quantity by $Q^{k,m}$, i.e., we expand $Q^{k}$ as
$Q^{k}=\sum_{m=1}^{k}Q^{k,m}.$ (13)
We first restrict the possible form of $k$-support operator $Q^{k,k}$. We here introduce the doubling-product operators $\mathit{\Delta}^{k}$ which is recursively defined as
$\mathit{\Delta}^{2}:=h,\hskip 10.0pt\mathit{\Delta}^{k}_{1,2,\ldots,k}:=[\mathit{\Delta}^{k-1}_{1,2,\ldots,k-1},h_{k-1,k}].$ (14)
It is shown in Ref. Hok25 that $Q^{k,k}$ should be proportional to $\mathit{\Delta}^{k}$ if $f(A)=[I_{1}\otimes A_{2},h_{12}]$ is injective and that the equation for 2-support operator $X$
$[X_{i,i+1},h_{i+1,i+2}]+[X_{i+1,i+2},h_{i,i+1}]=0$ (15)
admits only solutions for $X$ proportional to $h$. We provide the proof outline of satisfying these two conditions below. Their full proofs are shown in the Supplemental material SM .
The former injective condition is confirmed as follows: Using the projection operator formula in the character theory, we show that if $f(A)=0$, then $f(A^{l})=0$ for all $l$ where $A^{l}$ is the projection of $A$ onto the space with the azimuthal quantum number $l$. Note that $A^{l}$ is expanded as $A^{l}=\sum_{m}k_{m}Y^{l,m}$ with coefficient $k_{m}$. However, for any $l$, $m$ and $l^{\prime}$, there exists $m^{\prime}$ such that $[Y^{l,m},Y^{l^{\prime},m^{\prime}}]\neq 0$. Comparing Eq. (9), we conclude that $f(A^{l})=0$ happens only if $A^{l}$ is trivial (i.e., proportional to the identity operator).
We next treat the latter condition. We expand $X$ as
$X=\sum_{l,l^{\prime},m,m^{\prime}}q_{l,m,l^{\prime},m^{\prime}}Y^{l,m}_{i}Y^{l^{\prime},m^{\prime}}_{i+1}$ (16)
and specify coefficients $q_{l,m,l^{\prime},m^{\prime}}$ step by step. Let $L:=\{l|c_{l}\neq 0\}$ be the set of integers $l$ with nonzero coefficient $c_{l}$ in Eq. (9). It is easy to show that a term $Y^{l,m}Y^{l^{\prime},m^{\prime}}$ appears in $X$ only if $l,l^{\prime}\in L$. A key observation is that $Y^{l,m}Y^{l^{\prime},m^{\prime}}$ with a nonzero coefficient should take the form of $Y^{l,m}Y^{l,-m}$. To prove this, we employ a lemma that the coefficient of $Y^{l,m}Y^{l^{\prime},m^{\prime}}$ and that of $Y^{l,m\pm 1}Y^{l^{\prime},m^{\prime}\mp 1}$ are linearly connected:
$q_{l,m,l^{\prime},m^{\prime}}\propto q_{l,m\pm 1,l^{\prime},m^{\prime}\mp 1}.$ (17)
Applying this relation repeatedly, only one of $m$ or $m^{\prime}$ finally exceeds $l$ or $l^{\prime}$ if $(l^{\prime},m^{\prime})\neq(l,-m)$, which implies zero coefficient. Since all the coefficients in this sequence are linearly connected, we arrive at $q_{l,m,l^{\prime},m^{\prime}}=0$ if $(l^{\prime},m^{\prime})\neq(l,-m)$. We finally confirm that all the coefficients of $Y^{l,m}Y^{l,-m}$ are proportional to those of $h$, which is the desired relation.
We then show that $Q^{k,k}$ is in fact zero. A key fact to show this claim is that linear relations in $[Q^{k},H]=0$ for $k$-support operators can be reduced to those in $[Q^{3},H]=0$ for 3-support operators Hok25 . In other words, if we show $Q^{3,3}=0$ only from the relation that all coefficients of 3-support operators in
$[Q^{3,3}_{i,i+1,i+2},h_{i,i+1}]+[Q^{3,3}_{i,i+1,i+2},h_{i+1,i+2}]$
$+[Q^{3,2}_{i,i+1},h_{i+1,i+2}]+[Q^{3,2}_{i+1,i+2},h_{i,i+1}]$ (18)
are zero, then we have proof of the absence of $k$-local conserved quantities with general $k$. By comparing the condition of our main theorem and the above relation, it suffices to show $Q^{3,2}=0$. Below, we show that this relation is a consequence of the $SU(2)$ symmetry and the time-reversal symmetry of the Hamiltonian (1).
First, since both $H$ and $Q^{3,3}$ have $SU(2)$ symmetry, the entire conserved quantity $Q^{3}$, and thus $Q^{3,2}$, should also have $SU(2)$ symmetry. An $SU(2)$-symmetric 2-support operator is expressed as in the form of the right-hand-side of Eq. (1), or equivalently Eq. (9).
On the other hand, since $H$ has time-reversal symmetry, the conserved quantity $Q^{3}$ is either time-symmetric or time-antisymmetric. These two cases correspond to the cases that the sum of all azimuthal quantum numbers in the operator is even (time-symmetric) and odd (time-antisymmetric). Since $Q^{3,3}$ is equal to the doubling-product operator $\mathit{\Delta}^{3}$, which is time-antisymmetric, $Q^{3}$, and thus $Q^{3,2}$, should also be time-antisymmetric. However, any $SU(2)$-symmetric 2-support operator (the right-hand-side of Eq. (9)) is time-symmetric, and thus $Q^{3,2}$ should be zero, which completes the proof.
Discussion.— We have proven that a single quantity $D$ given in Eq. (2) sharply distinguishes two possible scenarios of isotropic spin chains: If $D$ has a 3-support operator, then the spin chain is non-integrable, i.e., no nontrivial local conserved quantity. If $D$ has no 3-support operator, then the spin chain satisfies the lowest-order Yang-Baxter equation (the Reshetikin condition) and has 3-local and 4-local conserved quantities in the form of Eqs. (3) and (4). Our result solves the Grabowski-Mathieu conjecture GM95-2 in the affirmative for isotropic spin chains, which is the first proof of the Grabowski-Mathieu conjecture applicable to systems whose integrability has still been unknown. Our result also serves as the classification theorem of integrability and non-integrability for the case with less than or equal to $S=13.5$.
In our theorem, a single relation serves as a criterion to distinguish the lowest-order Yang-Baxter equation and non-integrability, which is an advantage compared to previously proposed criteria GM95-2 ; Hok25 which contain search problems of operator equations. Provided conjecture 2 by Grabowski and Mathieu GM95-2 , the integrability or non-integrability of isotropic spin chains is completely determined by the same single relation. Such a simple criterion of integrability and non-integrability for a wider class of spin systems beyond isotropic spin chains is worth investigating.
Our result shows that all integrable isotropic spin chains satisfy the lowest-order Yang-Baxter equation. Moreover, restricting isotropic spin chains with $S\leq 13.5$, all integrable isotropic spin chains satisfy the Yang-Baxter equation. This shows high contrast to the Hubbard model, which is solvable but does neither satisfy the lowest-order Yang-Baxter equation nor have a boost operator generating local conserved quantities. On the other hand, the Hubbard model admits a generalized version of the Yang-Baxter equation and a generalized boost operator Lin01 ; Lee21 . On the basis of the above, our theorem based on the non-integrability proof sheds new light on the strongness and the limitation of the Yang-Baxter equation from the perspective of non-integrability, implying that the Yang-Baxter equation is the only solution method of integrable systems in some classes. We note that the technique of non-integrability proofs is also extended to write down local conserved quantities of integrable systems NF20 ; YF23 . The non-integrability proof approach will pave the way to clarify how universally the Yang-Baxter equation applies, which merits further research.
Acknowledgments. — It is a pleasure to thank Yuki Furukawa, for his essential contribution to Lemma. S.2 on the character theory. We also thank Akihiro Hokkyo, Yuuya Chiba, and Hal Tasaki for the fruitful discussion. This work is supported by JST ERATO Grant No. JPMJER2302, Japan.
Supplemental Material for
“Dichotomy theorem distinguishing non-integrability and the lowest-order Yang-Baxter equation for isotropic spin chains”
Naoto Shiraishi1 and Mizuki Yamaguchi1
1Graduate school of arts and sciences, The University of Tokyo
In this Supplemental Material, we first present brief reviews of integrability and non-integrability, respectively. These reviews will serve as a short summary of the research field of non-integrability, as well as a means to place our results properly in this context. We then provide proofs of several lemmas used in the main text.
I Brief review of techniques and results on integrable systems
I.1 Yang-Baxter equation and the quantum inverse scattering method
We first explain the Yang-Baxter equation (quantum inverse scattering method), which is a standard tool to solve the integrable systems and derive a family of infinite local conserved quantities. A detailed explanation can be seen in e.g., a lecture note by Fadeev s-Fad . Consider a one-parameter family of matrix $R_{ij}(\lambda)$ on two sites $i$ and $j$ with spin $S$ such that $R(0)$ is the swap operator $\Pi$, which is called $R$ matrix. The Yang-Baxter equation is then expressed as
$R_{12}(\lambda)R_{23}(\lambda+\mu)R_{23}(\mu)=R_{23}(\mu)R_{13}(\lambda+\mu)R_{12}(\lambda).$ (S.1)
If there exists $R$ matrix satisfying the Yang-Baxter equation, then we can construct an integrable Hamiltonian by the following procedure. We consider $N$ site $1,\ldots,N$ and one auxiliary site $a$ all of which have spin $S$. Using $R$ matrix, we introduce the monodromy matrix $M(\lambda)$ on sites $1,2,\ldots,N$ and $a$, and the transfer matrix $T(\lambda)$ on sites $1,2,\ldots,N$ as
$M(\lambda)$ $:=\prod_{i=1}^{N}R_{ia}(\lambda),$ (S.2)
$T(\lambda)$ $:=\mathrm{Tr}_{a}[M(\lambda)].$ (S.3)
Thanks to the Yang-Baxter equation, two transfer matrices commute
$[T(\lambda),T(\lambda^{\prime})]$ $=0,$ (S.4)
which can be demonstrated as follows. Since two operators acting on different sites commute, what we need to show is the middle equality of
$T(\lambda)T(\lambda^{\prime})=\mathrm{Tr}_{a,b}[\prod_{i}R_{ia}(\lambda)R_{ib}(\lambda^{\prime})]=\mathrm{Tr}_{a,b}[\prod_{i}R_{ib}(\lambda^{\prime})R_{ia}(\lambda)]=T(\lambda^{\prime})T(\lambda).$ (S.5)
Plugging $I=R^{-1}_{ab}(\lambda+\lambda^{\prime})R_{ab}(\lambda+\lambda^{\prime})$ to the left end of the product and applying the Yang-Baxter equation (S.1) repeatedly, we have
$\mathrm{Tr}_{a,b}[\prod_{i}R_{ia}(\lambda)R_{ib}(\lambda^{\prime})]$
$=$ $\mathrm{Tr}_{a,b}[R^{-1}_{ab}(\lambda+\lambda^{\prime})R_{ab}(\lambda+\lambda^{\prime})R_{1a}(\lambda)R_{1b}(\lambda^{\prime})R_{2a}(\lambda)R_{2b}(\lambda^{\prime})\cdots R_{Na}(\lambda)R_{Nb}(\lambda^{\prime})]$
$=$ $\mathrm{Tr}_{a,b}[R^{-1}_{ab}(\lambda+\lambda^{\prime})R_{1b}(\lambda^{\prime})R_{1a}(\lambda)R_{ab}(\lambda+\lambda^{\prime})R_{2a}(\lambda)R_{2b}(\lambda^{\prime})\cdots R_{Na}(\lambda)R_{Nb}(\lambda^{\prime})]$
$=$ $\mathrm{Tr}_{a,b}[R^{-1}_{ab}(\lambda+\lambda^{\prime})R_{1b}(\lambda^{\prime})R_{1a}(\lambda)R_{2b}(\lambda^{\prime})R_{2a}(\lambda)R_{ab}(\lambda+\lambda^{\prime})\cdots R_{Na}(\lambda)R_{Nb}(\lambda^{\prime})]$
$=$ $\cdots$
$=$ $\mathrm{Tr}_{a,b}[R^{-1}_{ab}(\lambda+\lambda^{\prime})R_{1b}(\lambda^{\prime})R_{1a}(\lambda)R_{2b}(\lambda^{\prime})R_{2a}(\lambda)\cdots R_{Nb}(\lambda^{\prime})R_{Na}(\lambda)R_{ab}(\lambda+\lambda^{\prime})]$
$=$ $\mathrm{Tr}_{a,b}[\prod_{i}R_{ib}(\lambda^{\prime})R_{ia}(\lambda)],$ (S.6)
which implies Eq. (S.4).
Consider the Taylor series of $f(T(\lambda))$ for any function $f$ with respect to $\lambda$ expanded as
$f(T(\lambda))=\sum_{n=0}^{\infty}F_{n}\lambda^{n}.$ (S.7)
Since Eq. (S.4) implies
$[f(T(\lambda),f(T(\mu))]=0,$ (S.8)
comparing the coefficient of $\lambda^{n}\mu^{m}$ in Eq. (S.8) we find
$[F_{n},F_{m}]=0.$ (S.9)
In particular, we set $f(x)=\ln x$ in the quantum inverse scattering method. Then, $F_{n-1}$ serves as the $n$-local conserved quantity, which is computed simply by differentiating $\ln T(\lambda)$ as
$Q^{n}=\frac{F_{n-1}}{(n-1)!}=\left.\frac{d^{n-1}}{d\lambda^{n-1}}\ln T(\lambda)\right|_{\lambda=0}.$ (S.10)
By regarding $Q^{2}$ as the Hamiltonian $H$, the obtained $Q^{n}$ is the $n$-local conserved quantity of $H$. The $n$-localness of $Q^{n}$ is confirmed by the boost operator argument presented in Sec. I.3
I.2 Reshetikin condition
Consider an $R$ matrix satisfying the Yang-Baxter equation (S.1). We can expand the $R$ matrix as
$R(\lambda)=\Pi+\lambda\Pi h+\lambda^{2}\Pi R^{(2)}+\lambda^{3}\Pi R^{(3)}+\cdots,$ (S.11)
where $h$ is the local Hamiltonian of $H$ and $\Pi$ is a swap operator.
For our purpose, it is useful to define
$R^{\prime}(\lambda)=\Pi R(\lambda),$ (S.12)
which is expanded as
$R^{\prime}(\lambda)=I+\lambda h+\lambda^{2}R^{(2)}+\lambda^{3}R^{(3)}+\cdots$ (S.13)
and satisfies a modulated version of the Yang-Baxter equation
$R^{\prime}_{12}(\lambda)R^{\prime}_{23}(\lambda+\mu)R^{\prime}_{12}(\mu)=R^{\prime}_{23}(\mu)R^{\prime}_{12}(\lambda+\mu)R^{\prime}_{23}(\lambda).$ (S.14)
We plug Eq. (S.13) into the Yang-Baxter equation (S.14) and compare the coefficients of $\lambda$ and $\mu$.
The coefficients of $\lambda\mu$ in Eq. (S.14) reads
$2R^{(2)}_{23}-h^{2}_{23}=2R^{(2)}_{12}-h^{2}_{12},$ (S.15)
which implies
$R^{(2)}=\frac{1}{2}h^{2}.$ (S.16)
The coefficients of $\lambda\mu^{2}$ reads
$R^{(3)}_{12}-R^{(3)}_{23}+\frac{1}{6}[h_{12}+h_{23},[h_{12},h_{23}]]-\frac{1}{6}h^{3}_{12}+\frac{1}{6}h^{3}_{23}=0.$ (S.17)
Defining unknown 2-support operator $E:=h^{3}-6R^{(3)}$, we arrive at the Reshetikin condition s-KS82 ; s-Ken92
$D_{123}=-[h_{12}+h_{23},[h_{12},h_{23}]]=E_{23}-E_{12}.$ (S.18)
I.3 Boost operator
We define the boost operator $B$ as
$B=\sum_{j}jh_{j,j+1}.$ (S.19)
We do not care about the effect of boundaries since it is irrelevant to our analysis.
We first suppose that a Hamiltonian satisfies the Yang-Baxter equation. Then, the corresponding transfer matrix $T(\lambda)$ satisfies
$[B,T(\lambda)]=\frac{d}{d\lambda}T(\lambda)+O(\lambda^{2}),$ (S.20)
which is shown as follows. Let ${\cal S}=T(0)$ be the shift operator by one site. We expand $T(\lambda+\varepsilon)$ up to both $\lambda$ and $\varepsilon$ at the first order as
$T(\lambda+\varepsilon)={\cal S}\left(1+H(\lambda+\varepsilon)+\sum_{j=1}^{N}h_{j,j+1}^{2}\lambda\varepsilon+2\sum_{i<j}h_{i,i+1}h_{j,j+1}\lambda\varepsilon+O(\lambda^{2},\varepsilon^{2})\right),$ (S.21)
which implies
$\frac{d}{d\lambda}T(\lambda)={\cal S}\left(H+\sum_{j=1}^{N}h_{j,j+1}^{2}\lambda+2\sum_{i<j}h_{i,i+1}h_{j,j+1}\lambda\right)+O(\lambda^{2}).$ (S.22)
On the other hand, $[B,T(\lambda)]$ up to the first order of $\lambda$ is calculated as
$[B,T(\lambda)]=$ $\sum_{j=1}^{N}[jh_{j,j+1},{\cal S}]+\sum_{i,j=1}^{N}[jh_{j,j+1},{\cal S}h_{i,i+1}]\lambda+O(\lambda^{2})$
$=$ ${\cal S}\left(H+\sum_{j=1}^{N}h_{j,j+1}^{2}\lambda+2\sum_{i<j}h_{i,i+1}h_{j,j+1}\lambda\right)+O(\lambda^{2}),$ (S.23)
which confirms Eq. (S.20).
Using Eq. (S.20), we find
$[B,T^{m}(\lambda)]=\sum_{l=0}^{m}T^{l}(\lambda)[B,T(\lambda)]T^{l-m-1}(\lambda)=\sum_{l=0}^{m}T^{l}(\lambda)\frac{d}{d\lambda}T(\lambda)T^{l-m-1}(\lambda)+O(\lambda^{2})=\frac{d}{d\lambda}T^{m}(\lambda)+O(\lambda^{2}),$ (S.24)
and thus, considering the Taylor expansion we find
$[B,f(T(\lambda))]=\frac{d}{d\lambda}f(T(\lambda))+O(\lambda^{2})$ (S.25)
for any analytic function $f$. Combining Eq. (S.25) with Eq. (S.10) by setting $f(x)=\ln x$, we find the recursive generation of local conserved quantities:
$[B,Q^{m}]=\left.\frac{d^{m}}{d\lambda^{m}}[B,\ln T(\lambda)]\right|_{\lambda=0}=\left.\frac{d^{m+1}}{d\lambda^{m+1}}\ln T(\lambda)\right|_{\lambda=0}=Q^{m+1}.$ (S.26)
This relation guarantees that $Q^{m}$ is indeed an $m$-local conserved quantity.
We next consider a general Hamiltonian $H$ on which the existence of the Yang-Baxter equation is not assumed. In this case, the obtained $m$-support operator $Q^{m}$ by Eq. (S.26) is no longer proven to be a conserved quantity.
The boost relation on the first order is
$Q^{3}=[B,H]=\sum_{i}[h_{i,i+1},h_{i+1,i+2}].$ (S.27)
Whether this $Q^{3}$ conserves or not is open at present. The conservation of $Q^{3}$ is expressed as
$[[B,H],H]=0,$ (S.28)
which is also equivalent to the Reshetikin condition (S.18).
If Eq. (S.28) (or the Reshetikin condition) holds, then $Q^{4}$ is also the 4-local conserved quantity, which is proven as
$[Q^{4},H]=[[B,Q^{3}],H]=-[[Q^{3},H],B]-[[H,B],Q^{3}]=0.$ (S.29)
On the other hand, whether $Q^{m}$ with $m\geq 5$ are conserved quantities or not has not yet been proven. Grabowski and Mathieu conjectured as conjecture 2 in Ref. s-GM95-2 that if Eq. (S.28) holds, then this boost operator indeed gives an infinite sequence of local conserved quantities.
II Brief review of non-integrability proofs
II.1 Several definitions and proof idea
We here briefly review several results and ideas of the proof of non-integrability. We first introduce several key notions and then explain the proof ideas. A detailed pedagogical introduction of non-integrability proofs can be found in Ref. s-Shi24 .
Throughout this Letter and Supplemental Material, we define non-integrability for a Hamiltonian with nearest-neighbor interaction by the absence of nontrivial local conserved quantities. To state it rigorously, we introduce the notion of a $k$-support operator such that its minimum contiguous support is among $k$ sites. For example, we state that $X_{4}Y_{5}Z_{6}$ is a 3-support operator and $X_{2}Y_{5}$ is a 4-support operator, since their contiguous supports are $\{4,5,6\}$ and $\{2,3,4,5\}$, respectively. A conserved quantity is a $k$-local conserved quantity if this quantity is given by a sum of up to $k$-support operators and cannot be expressed by a sum of at most $k-1$-support operators. The absence of nontrivial local conserved quantities in this case is defined by the absence of $k$-local conserved quantities for $k=O(1)$ with $k\geq 3$. Here, the word “nontrivial” means that the size of its contiguous support is strictly larger than that of the Hamiltonian.
The proof of non-integrability of spin chains is performed in the following procedure. We first expand the operator space acting on a single spin $S$ with respect to a proper basis. Then, we expand a candidate of a shift-invariant local conserved quantity up to $k$-support operators as
$Q^{k}=\sum_{l=1}^{k}Q^{k,l},\hskip 15.0ptQ^{k,l}:=\sum_{\bm{A}^{l}}\sum_{i=1}^{L}q_{\bm{A}^{l}}\bm{A}^{l}_{i}$ (S.30)
with coefficients $q_{\bm{A}^{l}}\in\mathbb{R}$. The sum of $\bm{A}^{l}$ runs over all possible sequences of basis operators with length $l$. We notice that the commutator of a $k$-support operator $Q^{k}$ and $H$ is an at most $k+1$-support operator, which guarantees the expansion as
$[Q^{k},H]=\sum_{l=1}^{k+1}\sum_{\bm{B}^{l}}\sum_{i=1}^{L}r_{\bm{B}^{l}}\bm{B}^{l}_{i}.$ (S.31)
Since the left-hand side can be expressed as a linear sum of $q_{\bm{A}}$ by inserting Eq. (S.30), and the conservation of $Q^{k}$ implies $r_{\bm{B}^{l}}=0$ for any $\bm{B}^{l}$, by comparing both sides of Eq. (S.31) we obtain many constraints (linear relations) on $q_{\bm{A}}$. Our goal is to show that these linear relations do not have solutions except for $q_{\bm{A}^{k}}=0$ for all $\bm{A}^{k}$, which means that $Q^{k}$ cannot be a $k$-local conserved quantity.
Concretely, we adopt the following two-step analysis.
Step 1: We analyze linear relations in Eq. (S.31) with $l=k+1$ and restrict $k$-support operators which may have nonzero coefficients to a specific form. In addition, we also show that the coefficients of the remaining $k$-support operators are linearly connected (i.e., $q_{\bm{A}}\propto q_{\bm{A}^{\prime}}$).
Step 2: We analyze linear relations in Eq. (S.31) with $l=k$ and demonstrate that one of the remaining coefficients of $k$-support operators is zero.
It is observed that most non-integrability proofs employ essentially the same proof technique as above Shi19 ; ST24 ; PL24-2 ; HYC24 ; YCS24-1 ; YCS24-2 ; FT25 , and Hokkyo Hok25 demonstrates that this proof technique indeed works if two operator equations respectively admit solutions of specific types.
In the following two subsections, we first explain the proof technique for non-integrability by employing $S=1/2$ XYZ chain with $z$ magnetic field
$H=\sum_{i}h^{(2)}_{i,i+1}+h^{(1)}_{i},\hskip 20.0pth^{(2)}_{i,i+1}=J_{X}X_{i}X_{i+1}+J_{Y}Y_{i}Y_{i+1}+J_{Z}Z_{i}Z_{i+1},\hskip 10.0pth^{(1)}_{i}=h_{Z}Z_{i}$ (S.32)
as an example, and then briefly describe its general form for general non-integrability proofs. Here, $X$, $Y$, and $Z$ are three Pauli operators. We emphasize that the $S=1/2$ XYZ chain with $z$ magnetic field is employed just for the explanation of the proof technique, and this spin chain is not isotropic (i.e., it is not an example of our main theorem). We also employ three Pauli matrices and the identity operator as the basis of operators ($2\times 2$ Hermitians), which is different from the noncommutative spherical harmonics $Y^{l,m}$.
II.2 Step 1 of non-integrability proofs
In step 1, we examine relations of $[Q^{k},H]=\sum_{l=1}^{k+1}\sum_{\bm{B}^{l}}\sum_{i=1}^{L}r_{\bm{B}^{l}}\bm{B}^{l}_{i}=0$ with $l=k+1$. To treat such relations transparently, we introduce a useful visualization of commutation relations. For example, for $k=4$, 5-support operator $\sum_{j}X_{j}Y_{j+1}X_{j+2}Y_{j+3}X_{j+4}$ in $[Q^{4},H]$ is generated by the following two commutators:
$-i[X_{j}Y_{j+1}X_{j+2}Z_{j+3},Y_{j+3}Y_{j+4}]$ $=-2X_{j}Y_{j+1}X_{j+2}X_{j+3}Y_{j+4},$ (S.33)
$-i[Z_{j+1}X_{j+2}X_{j+3}Y_{j+4},X_{j}X_{j+1}]$ $=2X_{j}Y_{j+1}X_{j+2}X_{j+3}Y_{j+4},$ (S.34)
where we dropped $\sum_{j}$, the summation over $j$, for visibility. Using the column expression, we express that $X_{j}Y_{j+1}X_{j+2}Y_{j+3}X_{j+4}$ is generated by these two commutation relations as follows:
$\begin{array}[]{rccccc}&X&Y&X&Z&\\ &&&&Y&Y\\ \hline\cr&X&Y&X&X&Y\end{array}\ \ \ \ \ \begin{array}[]{rccccc}&&Z&X&X&Y\\ &X&X&&&\\ \hline\cr&X&Y&X&X&Y\end{array},$
where we dropped all the coefficients in the outputs for brevity. The horizontal positions represent the spatial positions of spin operators. The column expression clearly shows that the operator $XYXXY$ in $[Q,H]$ is generated only by the above two commutators. Recalling the condition $r_{XYXXY}=0$, we have the following relation
$-q_{XYXZ}+q_{ZXXY}=0.$ (S.35)
Most of $k+1$-support operators are generated by two commutators. On the other hand, some $k+1$-support operators are generated only by one commutator. An example in $k=4$ is $YYYXZ$, which is generated only by
$\begin{array}[]{rccccc}&Y&Y&Y&Z&\\ &&&&Y&Y\\ \hline\cr&Y&Y&Y&X&Y\end{array}.$ (S.36)
Here, the other commutator vanishes since no 4-support operator satisfies the following relation:
$\ &Y&Y&&&\\ \hline\cr&Y&Y&Y&X&Y\end{array}.$ (S.37)
Then, Eq. (S.36) leads to
$-J_{Y}q_{YYYZ}=0,$ (S.38)
which directly implies $q_{YYYZ}=0$.
In addition, by considering commutators generating $XZYYZ$ as
$\begin{array}[]{ccccc}&Y&Y&Y&Z\\ X&X&&&\\ \hline\cr X&Z&Y&Y&Z\end{array}\hskip 15.0pt\begin{array}[]{ccccc}X&Z&Y&X&\\ &&&Z&Z\\ \hline\cr X&Z&Y&Y&Z\end{array},$ (S.39)
we readily have
$-J_{X}q_{YYYZ}-J_{Z}q_{XZYX}=0.$ (S.40)
However, since we have already shown $q_{YYYZ}=0$, we obtain $q_{XZYX}=0$.
Through similar arguments, we can show that a $k$-support operator has zero coefficient except for the case that this $k$-support operator can be constructed by repeatedly taking commutation relations with three interaction terms $XX$, $YY$, $ZZ$ with one-site shift as $[[[X_{i}X_{i+1},Y_{i+1}Y_{i+2}],X_{i+2}X_{i+3}],Z_{i+3}Z_{i+4}]$. From a more general perspective, such a $k$-support operator is an operator appearing in the doubling-product operators $\mathit{\Delta}^{k}$ which is recursively defined as
$\mathit{\Delta}^{2}:=h^{(2)},\hskip 10.0pt\mathit{\Delta}^{k}_{1,2,\ldots,k}:=[\mathit{\Delta}^{k-1}_{1,2,\ldots,k-1},h^{(2)}_{k-1,k}].$ (S.41)
In addition, it is easy to confirm that all the remaining coefficients of $k$-support operators should be proportional to those of $\mathit{\Delta}^{k}$. In summary, we find
$Q^{k,k}=c\mathit{\Delta}^{k}$ (S.42)
with a constant $c$. This condition for $k=3$ means
$Q^{3,3}=c\left(J_{X}J_{Y}XZY+J_{Y}J_{Z}YXZ+J_{Z}J_{X}ZYX-J_{Y}J_{X}YZX-J_{Z}J_{Y}ZXY-J_{X}J_{Z}XYZ\right).$ (S.43)
This type of argument appears not only in the non-integrability proof for the $S=1/2$ XYZ chain with $z$ magnetic field but also in various non-integrability proofs s-Shi19 ; s-Shi24 ; s-ST24 ; s-PL24-2 ; s-HYC24 ; s-YCS24-1 ; s-YCS24-2 ; s-Shi25 ; s-FT25 . By a careful observation of these facts, Hokkyo s-Hok25 demonstrated the wide applicability of this type of argument deriving Eq. (S.42) if (i) $f(A)=[I_{1}\otimes A_{2},h^{(2)}_{12}]$ and $g(A)=[A_{1}\otimes I_{2},h^{(2)}_{12}]$ are injective, and (2) the operator equation for 2-support operator $X$
$[X_{i,i+1},h^{(2)}_{i+1,i+2}]+e^{2\pi i/l}[X_{i+1,i+2},h^{(2)}_{i,i+1}]=0$ (S.44)
only has a solution $X\propto h^{(2)}$, where $l$ is a proper integer which divides the system size $L$. The form of Eq. (S.42) naturally arises since $k+1$-support operators in $[Q,H]$ on sites $11,2,\ldots,k+$ 1 come only from
$[Q^{k,k}_{1,\ldots,k},h^{(2)}_{k,k+1}]+[Q^{k,k}_{2,\ldots,k+1},h^{(2)}_{1,2}].$ (S.45)
Imposing all $k+1$-support operators to have zero coefficients and comparing these two terms, we notice that $Q^{k,k}_{1,\ldots,k}$ should take the form of $[Y_{2,\ldots,k},{h^{(2)}}_{1,2}]$ with a $k-1$-support operator $Y$. Applying similar arguments repeatedly, we obtain the doubling-product operator (S.42) if the last step (Eq. (S.44)) holds.
II.3 Step 2 of non-integrability proofs
In step 2, we examine $k$-support operators in $[Q,H]$ and demonstrate that all the remaining coefficients of $k$-support operators in $Q$ (i.e., $Q^{k,k}$) to be zero. Thanks to Eq. (S.44) showing that all coefficients share the same factor $c$, it suffices to show the coefficient of one of the remaining $k$-support operators in $Q$ to be zero.
All the $k$-support operators in $[Q,H]$ on sites $1,2,\ldots,k$ come only from
$[Q^{k,k-1}_{1,\ldots,k-1},h^{(2)}_{k-1,k}]+[Q^{k,k-1}_{2,\ldots,k},h^{(2)}_{1,2}]+\sum_{i=1}^{k-1}[Q^{k,k},h^{(2)}_{i,i+1}]+\sum_{i=1}^{k}[Q^{k,k},h^{(1)}_{i}].$ (S.46)
Note that this is a general expression and isotropic spin chains, which we treat in the main text, do not have the $h^{(1)}$ term. In these linear relations, $Q^{k,k}$ is known, while $Q^{k,k-1}$ is unknown. We extract several proper relations from these linear relations and solve them to show that one of the remaining coefficients of $k$-support operators to be zero.
We first treat the $S=1/2$ XYZ chain with $z$ magnetic field. Although our result holds for general $k$, we here demonstrate the proof only for $k=3,4,5$ in order to see what happens clearly. For $k=3$, we consider commutators generating $YYZ$ and $ZXX$ as
$\begin{array}[]{ccc}X&Y&Z\\ Z&&\\ \hline\cr Y&Y&Z\end{array}\hskip 15.0pt\begin{array}[]{ccc}Y&X&Z\\ &Z&\\ \hline\cr Y&Y&Z\end{array}\hskip 15.0pt\begin{array}[]{ccc}Y&X&\\ &Z&Z\\ \hline\cr Y&Y&Z\end{array}$ (S.47)
and
$\begin{array}[]{ccc}Z&X&Y\\ &&Z\\ \hline\cr Z&X&X\end{array}\hskip 15.0pt\begin{array}[]{ccc}Z&Y&X\\ &Z&\\ \hline\cr Z&X&X\end{array}\hskip 15.0pt\begin{array}[]{ccc}&Y&X\\ Z&Z&\\ \hline\cr Z&X&X\end{array},$ (S.48)
which lead to
$-h_{Z}q_{XYZ}-h_{Z}q_{YXZ}-J_{Z}q_{YX}$ $=0,$ (S.49)
$h_{Z}q_{ZXY}+h_{Z}q_{ZYX}+J_{Z}q_{YX}$ $=0.$ (S.50)
Summing these two relations and plugging relations on $q_{\bm{A}}$ shown in Eq. (S.43), we find
$-2\left(1-\frac{J_{Y}}{J_{X}}\right)h_{Z}q_{XYZ}=0.$ (S.51)
This relation implies that $q_{XYZ}=0$ as long as $h_{Z}\neq 0$ (XYZ model) and $J_{X}\neq J_{Y}$ (XXZ model with $z$ magnetic field).
For $k=4$, we consider commutators generating $YYYX$, $ZXYZ$, and $XYXX$ as
$\begin{array}[]{cccc}X&Y&Y&X\\ Z&&&\\ \hline\cr Y&Y&Y&X\end{array}\hskip 15.0pt\begin{array}[]{cccc}Y&X&Y&X\\ &Z&&\\ \hline\cr Y&Y&Y&X\end{array}\hskip 15.0pt\begin{array}[]{cccc}Y&Y&Z&\\ &&X&X\\ \hline\cr Y&Y&Y&X\end{array}$ (S.52)
and
$\begin{array}[]{cccc}Z&Y&Y&Z\\ &Z&&\\ \hline\cr Z&X&Y&Z\end{array}\hskip 15.0pt\begin{array}[]{cccc}Z&X&X&Z\\ &&Z&\\ \hline\cr Z&X&Y&Z\end{array}\hskip 15.0pt\begin{array}[]{cccc}&Y&Y&Z\\ Z&Z&&\\ \hline\cr Z&X&Y&Z\end{array}\hskip 15.0pt\begin{array}[]{cccc}Z&X&X&\\ &&Z&Z\\ \hline\cr Z&X&Y&Z\end{array}$ (S.53)
and
$\begin{array}[]{cccc}X&Y&X&Y\\ &&&Z\\ \hline\cr X&Y&X&X\end{array}\hskip 15.0pt\begin{array}[]{cccc}X&Y&Y&X\\ &&Z&\\ \hline\cr X&Y&X&X\end{array}\hskip 15.0pt\begin{array}[]{cccc}&Z&X&X\\ X&X&&\\ \hline\cr X&Y&X&X\end{array},$ (S.54)
which lead to
$-h_{Z}q_{XYYX}-h_{Z}q_{YXYX}-J_{X}q_{YYZ}$ $=0$ (S.55)
$h_{Z}q_{ZYYZ}-h_{Z}q_{ZXXZ}+J_{Z}q_{YYZ}-J_{Z}q_{ZXX}$ $=0$ (S.56)
$h_{Z}q_{XYXY}+h_{Z}q_{XYYX}+J_{X}q_{ZXX}$ $=0.$ (S.57)
Summing Eq. (S.55), Eq. (S.56) with multiplying $J_{X}/J_{Z}$, and Eq. (S.57), and combining Eq. (S.42) (which can be expanded as Eq. (S.43)), we obtain
$-3h_{Z}\left(1-\frac{J_{Y}}{J_{X}}\right)h_{Z}q_{XYYX}=0.$ (S.58)
This relation implies that $q_{XYYX}=0$ as long as $h_{Z}\neq 0$ (XYZ model) and $J_{X}\neq J_{Y}$ (XXZ model with $z$ magnetic field).
For $k=5$, we consider commutators $YYYZY$, $ZXYYX$, $XYXYZ$, and $YZYXX$ as
$\begin{array}[]{ccccc}X&Y&Y&Z&Y\\ Z&&&&\\ \hline\cr Y&Y&Y&Z&Y\end{array}\hskip 15.0pt\begin{array}[]{ccccc}Y&X&Y&Z&Y\\ &Z&&&\\ \hline\cr Y&Y&Y&Z&Y\end{array}\hskip 15.0pt\begin{array}[]{ccccc}Y&Y&Y&X&\\ &&&Y&Y\\ \hline\cr Y&Y&Y&Z&Y\end{array}$ (S.59)
and
$\begin{array}[]{ccccc}Z&Y&Y&Y&X\\ &Z&&&\\ \hline\cr Z&X&Y&Y&X\end{array}\hskip 15.0pt\begin{array}[]{ccccc}Z&X&X&Y&X\\ &&Z&\\ \hline\cr Z&X&Y&Y&X\end{array}\hskip 15.0pt\begin{array}[]{ccccc}&Y&Y&Y&X\\ Z&Z&&&\\ \hline\cr Z&X&Y&Y&X\end{array}\hskip 15.0pt\begin{array}[]{ccccc}Z&X&Y&Z&\\ &&&X&X\\ \hline\cr Z&X&Y&Y&X\end{array}$ (S.60)
and
$\begin{array}[]{ccccc}X&Y&Y&Y&Z\\ &&Z&&\\ \hline\cr X&Y&X&Y&Z\end{array}\hskip 15.0pt\begin{array}[]{ccccc}X&Y&X&X&Z\\ &&&Z&\\ \hline\cr X&Y&X&Y&Z\end{array}\hskip 15.0pt\begin{array}[]{ccccc}&Z&X&Y&Z\\ X&X&&&\\ \hline\cr X&Y&X&Y&Z\end{array}\hskip 15.0pt\begin{array}[]{ccccc}X&Y&X&X&\\ &&&Z&Z\\ \hline\cr X&Y&X&Y&Z\end{array}$ (S.61)
and
$\begin{array}[]{ccccc}Y&Z&Y&Y&X\\ &&&Z&\\ \hline\cr Y&Z&Y&X&X\end{array}\hskip 15.0pt\begin{array}[]{ccccc}Y&Z&Y&X&Y\\ &&&&Z\\ \hline\cr Y&Z&Y&X&X\end{array}\hskip 15.0pt\begin{array}[]{ccccc}&X&Y&X&X\\ Y&Y&&&\\ \hline\cr Y&Z&Y&X&X\end{array},$ (S.62)
which lead to
$-h_{Z}q_{XYYZY}-h_{Z}q_{YXYZY}+J_{Y}q_{YYYX}$ $=0$ (S.63)
$h_{Z}q_{ZYYYX}-h_{Z}q_{ZXXYX}+J_{Z}q_{YYYX}+J_{X}q_{ZXYZ}$ $=0$ (S.64)
$h_{Z}q_{XYYYZ}-h_{Z}q_{XYXXZ}+J_{X}q_{ZXYZ}-J_{Z}q_{XYXX}$ $=0$ (S.65)
$h_{Z}q_{YZYYX}+h_{Z}q_{YZYXY}+J_{Y}q_{XYXX}$ $=0.$ (S.66)
Summing Eq. (S.63), Eq. (S.64) with multiplying $-J_{Y}/J_{Z}$, Eq. (S.65) with multiplying $J_{Y}/J_{Z}$, and Eq. (S.66), and combining Eq. (S.42), we obtain
$-4h_{Z}\left(1-\frac{J_{Y}}{J_{X}}\right)h_{Z}q_{XYYZY}=0.$ (S.67)
This relation implies that $q_{XYYZY}=0$ as long as $h_{Z}\neq 0$ (XYZ model) and $J_{X}\neq J_{Y}$ (XXZ model with $z$ magnetic field).
As seen from the above $k=3$, 4, and 5 cases, the structure of step 2 is very similar among all $k$. In fact, the arguments of general $k$ can be obtained by extending those for $k=3$, specifically by elongating the operators appearing in commutation relations for $k=3$ through the alternating insertion of $XX$ and $YY$, in order to construct $k$-support and $k-1$-support operators from 3-support and 2-support operators. Hence, if we succeed in proving that one of the remaining 3-support operators has zero coefficients, then we can extend this argument to general $k$.
We encounter similar situation in various non-integrability proofs s-Shi19 ; s-ST24 ; s-PL24-2 ; s-HYC24 ; s-YCS24-1 ; s-YCS24-2 ; s-FT25 . Observing this situation, Hokkyo s-Hok25 demonstrated that this elongation/reduction between general $k$-support operators and that for $k=3$ indeed works if (i) Eq. (S.42) holds and (ii) we succeed in proving the absence of 3-local conserved quantity by step 2 (i.e., we need not examine 2-support and 1-support operators in $[Q,H]$).
III Derivation of injective condition
Here we show that $f(A)=[I_{1}\otimes A_{2},h_{12}]$ is injective for any nontrivial isotropic $h_{12}$. Recalling the expression (9) and the fact that $\{Y^{l,m}\}_{l,m}$ is a basis of the operator space, our claim is equivalent to the following lemma.
Lemma S.1.
Fix the azimuthal quantum number $l>0$. If $[A,Y^{l,m}]=0$ holds for all $-l\leq m\leq l$, then $A$ is proportional to the identity operator $I$.
We prove this lemma by two steps. Below, we denote by $V^{l}={\rm span}\{Y^{l,m}\}$ an operator subspace with azimuthal quantum number $l$.
Lemma S.2.
Fix the azimuthal quantum number $l>0$. If $[A,Y^{l,m}]=0$ holds for all $-l\leq m\leq l$, then $[A^{l},Y^{l,m}]=0$ holds for all $-l\leq m\leq l$, where $A^{l}$ is the projection of $A$ onto $V^{l}$.
We learned the following proof by Yuki Furukawa.
Proof.
[…]
Lemma S.3.
For any $l_{1},l_{2}>0$ and $m_{1}$, there exists $m_{2}$ such that $[Y^{l_{1},m_{1}},Y^{l_{2},m_{2}}]\neq 0$.
Proof.
We set $m_{2}=-l_{2}$ for $l_{1}\geq 0$ and $m_{2}=l_{2}$ for $l_{1}<0$. If $l_{1}+l_{2}-1\leq 2S$, we set $l=l_{1}+l_{2}-1$, and if $l_{1}+l_{2}-1>2S$ we set $l$ as one of $2S$ or $12S-$ 1 such that $l_{1}+l_{2}+l$ is odd. With this choice, $m=m_{1}+m_{2}$ satisfies $\left|m\right|\leq\left|l_{2}\right|\leq l$.
Under the above choice, $[Y^{l_{1},m_{1}},Y^{l_{2},m_{2}}]$ has $Y^{l,m}$ with a nonzero coefficient. The non-vanishing of the coefficient (i.e., no accidental zero) is confirmed in Lemma. S.12. ∎
Combining these two lemmas, we find that $[A,Y^{l,m}]=0$ for all $-l\leq m\leq l$ implies $A\propto I$, which completes the proof of Lemma. S.1.
IV Derivation of Eq.(9)
Here we shall show that an SU(2)-symmetric spin chain Hamiltonian $H=\sum_{i}h_{i,i+1}$ can be expressed as Eq. (9), which we recast below:
$h_{i,i+1}=\sum_{l=0}^{2S}c_{l}\left(\sum_{m=-l}^{l}Y^{l,m}_{i}Y^{l,-m}_{i+1}\right).$ (S.68)
Note that the SU(2) symmetry means
$[h_{i,i+1},S^{0}]$ $=0$ (S.69)
$[h_{i,i+1},S^{\pm 1}]$ $=0$ (S.70)
with
$S^{m}$ $:=\sum_{i=1}^{N}Y_{i}^{1,m}\hskip 10.0pt(m=0,\pm 1).$ (S.71)
Below we start with a general expression of $h_{i,i+1}$ without any symmetry
$h_{i,i+1}=\sum_{l,l^{\prime}=0}^{2S}\sum_{m=-l}^{l}\sum_{m^{\prime}=-l^{\prime}}^{l^{\prime}}c_{lml^{\prime}m^{\prime}}Y^{l,m}_{i}Y^{l^{\prime},m^{\prime}}_{i+1}$ (S.72)
and restrict its possible form to Eq. (S.68) by imposing the SU(2) symmetry: (S.69) and (S.70).
Lemma S.4.
The coefficient $c_{l,l^{\prime},m,m^{\prime}}$ in Eq. (S.72) for a spin chain with Eq. (S.69) is zero if $m\neq-m^{\prime}$.
Proof.
Since the total magnetic quantum number is the eigenvalue of the commutator with $S^{0}$, i.e.,
$[Y_{i}^{l,m},S^{0}]=mY_{i}^{l,m},$ (S.73)
we find
$[h_{i,i+1},S^{0}]=\sum_{l,l^{\prime},m,m^{\prime}}c_{lml^{\prime}m^{\prime}}[Y_{i}^{l,m}Y_{i+1}^{l^{\prime},m^{\prime}},S^{0}]=\sum_{l,l^{\prime},m,m^{\prime}}c_{lml^{\prime}m^{\prime}}(m+m^{\prime})Y_{i}^{l,m}Y_{i+1}^{l^{\prime},m^{\prime}}.$ (S.74)
Comparing this relation with Eq. (S.69), we find $c_{l,l^{\prime},m,m^{\prime}}=0$ for $m\neq-m^{\prime}$. ∎
Lemma S.5.
The coefficient $c_{l,l^{\prime},m,-m}$ in Eq. (S.72) for a spin chain with Eq. (S.69) and Eq. (S.70) is zero if $l\neq l^{\prime}$.
Proof.
Recall that the coefficient $K^{l_{1},l_{2},l}_{m_{1},m_{2}}$ in Eq. (8) (in the main text), which is proportional to the Clebsh-Gordon coefficient, is zero if the triangular condition $\left|l_{1}-l_{2}\right|\leq l\leq l_{1}+l_{2}$ is not satisfied. In addition, we also have $K^{l_{1},l_{2},l}_{m_{1},m_{2}}=K^{l_{2},l_{1},l}_{m_{2},m_{1}}$ if $l_{1}+l_{2}+l$ is even and $K^{l_{1},l_{2},l}_{m_{1},m_{2}}=-K^{l_{2},l_{1},l}_{m_{2},m_{1}}$ if $l_{1}+l_{2}+l$ is odd. Setting $l_{2}=1$, we find
$[Y_{i}^{l,m},S^{\pm 1}]=[Y_{i}^{l,m},Y_{i}^{1,\pm 1}]=2K^{l,1,l}_{m,\pm 1}Y_{i}^{l,m\pm 1}.$ (S.75)
The coefficient $K^{l,1,l}_{m,\pm 1}$ is always nonzero for $\left|m\right|\leq l$ and $\left|m\pm 1\right|\leq l$ (for its proof, see Lemma. S.11), and zero otherwise. Using this relation, we find
$[Y_{i}^{l,m}Y_{i+1}^{l^{\prime},-m},S^{\pm 1}]=2K^{l,1,l}_{m,\pm 1}Y_{i}^{l,m\pm 1}Y_{i+1}^{l^{\prime},-m}+2K^{l^{\prime},1,l^{\prime}}_{-m,\pm 1}Y_{i}^{l,m}Y_{i+1}^{l^{\prime},-m\pm 1}.$ (S.76)
We shall show that $c_{l,l^{\prime},m,-m}=0$ for $l\neq l^{\prime}$. Without loss of generality, we assume $m\geq 0$ and $l<l^{\prime}$. Other cases are treated in similar manners. We insert Eq. (S.76) for $S^{-1}$ into Eq. (S.70) and compare the coefficients of $Y_{i}^{l,m}Y_{i+1}^{l^{\prime},-m-1}$, which leads to a linear relation
$02K^{l^{\prime},1,l^{\prime}}_{-m,-1}c_{l,m,l^{\prime},-m}+2K^{l,1,l}_{m+1,-1}c_{l,m+1,l^{\prime},-m-1}=$ 0 . (S.77)
If $\left|m+1\right|\leq l$ (and hence $\left|-m-1\right|\leq l^{\prime}$) is satisfied, then both coefficients $K$ are nonzero and we have
$c_{l,m,l^{\prime},-m}\propto c_{l,m+1,l^{\prime},-m-1}.$ (S.78)
Thus, replacing $m$ by $m+n$ ($n=1,2,\ldots$) and applying Eq. (S.77) repeatedly, we find
$c_{l,m,l^{\prime},-m}\propto c_{l,m+1,l^{\prime},-m-1}\propto c_{l,m+2,l^{\prime},-m-2}\propto\cdots\propto c_{l,l,l^{\prime},-l}.$ (S.79)
On the other hand, at $c_{l,l,l^{\prime},-l}$ we have $K^{l,1,l}_{l+1,-1}=0$ but $K^{l^{\prime},1,l^{\prime}}_{-l,-1}\neq 0$, which leads to
$02K^{l^{\prime},1,l^{\prime}}_{-m,-1}c_{l,l,l^{\prime},-l}=$ 0 . (S.80)
Plugging this relation into Eq. (S.79), we arrive at $c_{l,m,l^{\prime},-m}=0$. ∎
Lemma S.6.
If $c_{l,m,l,-m}$ is nonzero with some $m$, then $c_{l,m^{\prime},l,-m^{\prime}}$ is also nonzero.
Proof.
A Similar argument to Eq. (S.77) for $S^{+1}$ yields
$c_{l,m,l^{\prime},-m}\propto c_{l,m-1,l^{\prime},-m+1}.$ (S.81)
Using this relation and Eq. (S.78), we find
$c_{l,l,l,-l}\propto c_{l,l-1,l,-l+1}\propto c_{l,l-2,l,-l+2}\propto\cdots\propto c_{l,-l,l,l},$ (S.82)
which directly implies the desired property. ∎
We renormalize $Y^{l,m}$ such that these constants of proportion are one, which leads to
$c_{l,l,l,-l}=c_{l,l-1,l,-l+1}=c_{l,l-2,l,-l+2}=\cdots=c_{l,-l,l,l}.$ (S.83)
By setting $c_{l}=c_{l,l,l,-l}$, we arrive at the desired expression (S.68).
V Demonstration that Eq.(15) has solution only $X\propto h$
We here show that any solution $X$ to Eq.(15) must be proportional to $h$. To show this, we first denote by $L:=\{l|c_{i}\neq 0\}$ the set of nonzero coefficients $c_{l}$ in Eq. (9). By assumption, $L\neq\emptyset$. We then introduce several lemmas.
Lemma S.7.
The coefficient of $Y^{l,m}Y^{l^{\prime},m^{\prime}}$ in $X$ with $l\notin L$ is zero.
Proof.
We expand $X$ as $X=\sum_{l,m}Y^{l,m}A^{l,m}$, where $A^{l,m}$ is generally not a single noncommutative spherical harmonics but their proper sum. We suppose $A^{l,m}$ is nontrivial for $l\notin L$ and derive a contradiction.
Lemma. S.1 guarantees that for any nontrivial $A^{l,m}$ there exist $l^{*}\in L$ and $m^{*}$ satisfying $[A^{l,m},Y^{l^{*},m^{*}}]\neq 0$. Then, an operator in the form of $Y^{l,m}_{i}[A^{l,m},Y^{l^{*},m^{*}}]_{i+1}Y^{l^{*},-m^{*}}_{i+2}$ is generated by $[X_{i,i+1},h_{i+1,i+2}]$ but is not generated by $[X_{i+1,i+2},h_{i,i+1}]$ since $l\notin L$. This means that $Y^{l,m}_{i}[A^{l,m},Y^{l^{*},m^{*}}]_{i+1}Y^{l^{*},-m^{*}}_{i+2}$ does not vanish in the left-hand side of Eq. (15), which is contradiction. ∎
Lemma S.8.
The coefficient of $Y^{l,m}Y^{l^{\prime},m^{\prime}}$ in $X$ with $l,l^{\prime}\in L$ has a zero coefficient if $l\neq l^{\prime}$ or $m\neq-m^{\prime}$.
Proof.
In this and the following proofs, we use the column expression of commutation relations introduced in Sec. II. Note that unlike Sec. II the last row in the column expression represents only one of the operators contained in this commutator, not all of them.
Consider commutators in $[X_{i,i+1},h_{i+1,i+2}]$ and $[X_{i+1,i+2},h_{i,i+1}]$ generating operators $Y^{l,m}_{i}Y^{1,\pm 1}_{i+1}Y^{l^{\prime},m^{\prime}\mp 1}_{i+2}$. The commutator $[X_{i,i+1},h_{i+1,i+2}]$ contains
$\begin{array}[]{ccc}Y^{l,m}&Y^{l^{\prime},m^{\prime}}&\\ &Y^{l^{\prime},-m^{\prime}\pm 1}&Y^{l^{\prime},m^{\prime}\mp 1}\\ \hline\cr Y^{l,m}&Y^{1,\pm 1}&Y^{l^{\prime},m^{\prime}\mp 1}\end{array}$ (S.84)
if $\left|m^{\prime}\right|\leq l^{\prime}$, and the commutator $[X_{i+1,i+2},h_{i,i+1}]$ contains
$\begin{array}[]{ccc}&Y^{l,m\pm 1}&Y^{l^{\prime},m^{\prime}\mp 1}\\ Y^{l,m}&Y^{l,-m}&\\ \hline\cr Y^{l,m}&Y^{1,\pm 1}&Y^{l^{\prime},m^{\prime}\mp 1}\end{array}$ (S.85)
if $\left|m\pm 1\right|\leq l$. These two relations imply a linear relation between coefficients of $Y^{l,m}Y^{l^{\prime},m^{\prime}}$ and $Y^{l,m\pm 1}Y^{l^{\prime},m^{\prime}\mp 1}$ as
$K^{l^{\prime},l^{\prime},1}_{m^{\prime},-m^{\prime}\pm 1}c_{l^{\prime}}q_{l,m,l^{\prime},m^{\prime}}+K^{l,l,1}_{m\pm 1,-m}c_{l}q_{l,m\pm 1,l^{\prime},m^{\prime}\mp 1}=0.$ (S.86)
Both coefficients $K^{l,l,1}_{m\pm 1,-m}$ and $K^{l^{\prime},l^{\prime},1}_{m^{\prime},-m^{\prime}\pm 1}$ are nonzero if $\left|m^{\prime}\right|\leq l^{\prime}$ and $\left|m\pm 1\right|\leq l$ (for its proof, see Lemma. S.10), which implies
$q_{l,m,l^{\prime},m^{\prime}}\propto q_{l,m\pm 1,l^{\prime},m^{\prime}\mp 1}.$ (S.87)
Applying Eq. (S.87) repeatedly, we obtain a sequence
$q_{l,m,l^{\prime},m^{\prime}}\propto q_{l,m\pm 1,l^{\prime},m^{\prime}\mp 1}\propto q_{l,m\pm 2,l^{\prime},m^{\prime}\mp 2}\propto q_{l,m\pm 3,l^{\prime},m^{\prime}\mp 3}\cdots.$ (S.88)
This sequence finally reaches $\left|m\right|=l$ or $\left|m^{\prime}\right|=l^{\prime}$.
On the other hand, if exactly one of $\left|m^{\prime}\right|\leq l^{\prime}$ or $\left|m\pm 1\right|\leq l$ is not satisfied, then the remaining coefficient is shown to be zero. In the case with $\left|m^{\prime}\right|\leq l^{\prime}$ and $\left|m\pm 1\right|>l$, we have $K^{l,l,1}_{m\pm 1,-m}=0$ and $K^{l^{\prime},l^{\prime},1}_{m^{\prime},-m^{\prime}\pm 1}\neq 0$, which leads to
$c_{l^{\prime}}q_{l,m,l^{\prime},m^{\prime}}=0.$ (S.89)
In the case with $\left|m^{\prime}\right|>l^{\prime}$ and $\left|m\pm 1\right|\leq l$, we have $K^{l^{\prime},l^{\prime},1}_{m^{\prime},-m^{\prime}\pm 1}=0$ and $K^{l,l,1}_{m\pm 1,-m}\neq 0$, which leads to
$c_{l}q_{l,m\pm 1,l^{\prime},m^{\prime}\mp 1}=0.$ (S.90)
Let us consider the case that one of $l=l^{\prime}$ or $m=-m^{\prime}$ is not satisfied. Due to the symmetry of $(l,m)\leftrightarrow(l^{\prime},m^{\prime})$ and $(m,m^{\prime})\leftrightarrow(-m,-m^{\prime})$, without loss of generality we can assume $l-m<l^{\prime}+m^{\prime}$. Other cases are treated in similar manners. In the case of $l-m<l^{\prime}+m^{\prime}$, the $q_{l,m+n,l^{\prime},m^{\prime}-n}$ sequence in Eq. (S.88) reaches $q_{l,l,l^{\prime},m^{\prime}+m-l}$. Noting $\left|m^{\prime}+m-l\right|\leq l^{\prime}$, we apply (S.89) and find that all the coefficients in this sequence including $q_{l,m,l^{\prime},m^{\prime}}$ are zero. ∎
Lemma S.9.
The coefficient of $Y^{l,m}Y^{l,-m}$ in $X$ is proportional to that in $h$.
Proof.
We have shown in Lemma. S.8 that $X$ claims
$X=\sum_{l,m}d_{l,m}Y^{l,m}Y^{l,-m},$ (S.91)
where we denote $d_{l,m}=q_{l,m,l,-m}$. This lemma is equivalent to
$\frac{d_{l,m}}{d_{l^{\prime},m^{\prime}}}=\frac{c_{l}}{c_{l^{\prime}}}$ (S.92)
for any $l,m,l^{\prime},m^{\prime}$.
We first consider the case of $l=l^{\prime}$. We employ the same sequence obtained in the proof of Lemma. S.8. We set $l=l^{\prime}$ and $m^{\prime}=-m$ in Eqs. (S.84) and (S.85), which leads to
$K^{l,l,1}_{-m,m\pm 1}c_{l}d_{l,m}+K^{l,l,1}_{m\pm 1,-m}c_{l}d_{l,m\pm 1}=0.$ (S.93)
Substituting $K^{l,l,1}_{a,b}=-K^{l,l,1}_{b,a}$, which follows from oddness of $l+l+1$, into the above relation and replacing $m$ by $m\pm n$ repeatedly, we arrive at
$d_{l,m}=d_{l,m\pm 1}=d_{l,m\pm 2}=\cdots,$ (S.94)
thereby confirming Eq. (S.92) in the case $l=l^{\prime}$. We denote the value of the above equation simply by $d_{l}$.
We then treat the case of $l\neq l^{\prime}$ and show $d_{l}/d_{l^{\prime}}=c_{l}/c_{l^{\prime}}$. Lemma. S.3 confirms that for any $l$ and $l^{\prime}$ there exist $m$ and $m^{\prime}$ such that $Y^{l,m}$ and $Y^{l^{\prime},m^{\prime}}$ do not commute. Let $A$ be an operator which has a nonzero coefficient $K_{a}$ in $[Y^{l,m},Y^{l^{\prime},m^{\prime}}]$. Then, we consider commutators in $[X_{i,i+1},h_{i+1,i+2}]$ and $[X_{i+1,i+2},h_{i,i+1}]$ generating operators $Y^{l,-m}_{i}A_{i+1}Y^{l^{\prime},-m^{\prime}}_{i+2}$. The commutator $[X_{i,i+1},h_{i+1,i+2}]$ contains
$\begin{array}[]{ccc}Y^{l,-m}&Y^{l,m}&\\ &Y^{l^{\prime},m^{\prime}}&Y^{l^{\prime},-m^{\prime}}\\ \hline\cr Y^{l,-m}&A&Y^{l^{\prime},-m^{\prime}}\end{array}$ (S.95)
and the commutator $[X_{i+1,i+2},h_{i,i+1}]$ contains
$\begin{array}[]{ccc}&Y^{l^{\prime},m^{\prime}}&Y^{l^{\prime},-m^{\prime}}\\ Y^{l,-m}&Y^{l,m}&\\ \hline\cr Y^{l,-m}&A&Y^{l^{\prime},-m^{\prime}}\end{array},$ (S.96)
which lead to
$K_{a}d_{l}c_{l^{\prime}}-K_{a}d_{l^{\prime}}c_{l}=0.$ (S.97)
This directly implies the desired relation $d_{l}/d_{l^{\prime}}=c_{l}/c_{l^{\prime}}$, which completes the proof. ∎
VI Absence of accidental zeros in some Clebsch-Gordan coefficients
In our analyses, we have taken commutation relations and derived nontrivial relations on coefficients. To perform these analyses, the coefficients of commutation relations $K$, which is proportional to the Clebsh-Gordan coefficient, should be nonzero. Almost all Clebsh-Gordan coefficients are nonzero if the triangle inequality raised in the main text is satisfied. However, it is known that there exist some accidental zeros in Clebsh-Gordan coefficients even if these two conditions are satisfied s-RR84 ; s-HHR09 ; s-Donbook . Although it is highly implausible, for completeness we here prove that the Clebsh-Gordan coefficients employed in our analyses are indeed nonzero, i.e., we do not face accidental zeros in our argument.
Lemma S.10.
The coefficient $K^{l_{1},l_{2},l}_{m_{1},m_{2}}$ proportional to the Clebsch-Gordan coefficient takes a nonzero value when $l_{1}=l_{2}$, $l=1$, and $m_{1}=-m_{2}\pm 1$.
Proof.
Thanks to the general expression of the Clebsch-Gordan coefficients s-Bohbook , $K^{l_{1},l_{2},l}_{m_{1},m_{2}}$ is proportional to
$\sum_{k=A}^{B}\frac{(-1)^{k}}{k!(l_{1}+l_{2}-l-k)!(l_{1}-m_{1}-k)!(l_{2}+m_{2}-k)!(l-l_{2}+m_{1}+k)!(l-l_{1}-m_{2}+k)!},$ (S.98)
where $A$ and $B$ are set to
$A$ $=\max(0,l_{2}-l-m_{1},l_{1}-l+m_{2}),$ (S.99)
$B$ $=\min(l_{1}+l_{2}-l,l_{1}-m_{1},l_{2}+m_{2}),$ (S.100)
with which all the entries of factorials are nonnegative. When $j=l_{1}=l_{2}$, $l=1$, and $m=m_{1}=-m_{2}\pm 1$, $A$ and $B$ are
$A$ $=\max(0,j-1-m,j-1-m\pm 1),$ (S.101)
$B$ $=\min(2j-1,j-m,j-m\pm 1),$ (S.102)
and thus $B=A$ in both cases of $\pm$. This means that Eq. (S.98) is a sum of a single nonzero summand, which is trivially nonzero. ∎
Lemma S.11.
The coefficient $K^{l_{1},l_{2},l}_{m_{1},m_{2}}$ proportional to the Clebsch-Gordan coefficient takes a nonzero value when $l_{1}=l$, $l_{2}=1$, and $m_{2}=\pm 1$.
Proof.
We again employ the expression (S.98). When $l_{1}=l$, $l_{2}=1$, and $m_{2}=\pm 1$, $A$ and $B$ are
$A$ $=\max(0,1-l-m_{1},-m_{2}),$ (S.103)
$B$ $=\min(1,l-m_{1},1-m_{2}).$ (S.104)
In the case of $m_{2}=1$, we have $A=B=0$, and in the case of $m_{2}=-1$, we have $A=B=1$. This means that Eq. (S.98) is a sum of a single nonzero summand, which is trivially nonzero. ∎
Lemma S.12.
The coefficient $K^{l_{1},l_{2},l}_{m_{1},m_{2}}$ proportional to the Clebsch-Gordan coefficient takes a nonzero value when $\left|m_{2}\right|=l_{2}$, $(l_{1},l_{2},l)$ satisfies the triangle relation, and $l_{1}+l_{2}+l$ is odd.
Proof.
We again employ the expression (S.98).
We first treat the case of $m_{2}=-l_{2}$. In this case, $A$ and $B$ are
$A$ $=\max(0,l_{2}-l-m_{1},l_{1}-l-m_{2}),$ (S.105)
$B$ $=\min(l_{1}+l_{2}-l,l_{1}-m_{1},0),$ (S.106)
which directly implies $A=B=0$.
We next treat the case of $m_{2}=l_{2}$. In this case, $A$ and $B$ are
$A$ $=\max(0,l_{2}-l-m_{1},l_{1}-l+l_{2}),$ (S.107)
$B$ $=\min(l_{1}+l_{2}-l,l_{1}-m_{1},2l_{2}),$ (S.108)
which directly implies $A=B=l_{1}+l_{2}-l$.
In both cases, Eq. (S.98) is a sum of a single nonzero summand, which is trivially nonzero. ∎